Air contains 80% of nitrogen no of moles of nitrogen resent in 56 litres of airat stp
 

Grade:8

1 Answers

Khimraj
5 years ago
volume of nitrogen = 56*4/5 = 44.8L
mole of nitrogen = 44.8/22.4 = 2 moles.
Hope it clears...........................................................